# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ 5.3 Deterministic Controllers
Finally, we analyze the case with a history-dependent deterministic controller and a Markov convex adversary. We consider the history-dependent policy such that $\pi_{2i}=a_{1},i=1,2,\ldots,$ if the controller sees state I1 and $\pi_{2i}=a_{2},i=1,2,\ldots,$ if the controller sees state I2. Then, for any Markov adversary $\kappa$, we always have
$v(I0,\pi,\kappa)=0>u^{*}(I0).$
Again, this suggests the absence of a DPP in this setting as well.
@@ 6 Connections with other Formulations
In this section, we elucidate the links between the formulations of DRMDP and the multistage stochastic programs (Shapiro et al., 2021, Chapter 3) within the realm of optimization literature. Additionally, we explore their connections with stochastic games, as introduced by Shapley (1953) in the economics literature.
@@ 6.1 Multistage Stochastic Programs
Early works on DRMDP, such as González-Trejo et al. (2002) and Nilim and El Ghaoui (2005), interpreted the max-min problem as dynamic interactions between a controller and an SA-rectangular adversary. Indeed, the presence of a Dynamic Programming Principle (DPP) in these scenarios implies that the interactions under optimal robust control are equivalent to another dynamic model wherein the controller and adversary sequentially make decisions. This interpretation results in what is now considered as a robustified variant of the multistage stochastic program (MSP) (Huang et al., 2017). In contrast, Nilim and El Ghaoui (2005)’s version of sequential max-min-max-min decisions (their Equation (5)) is positioned before time 0.
As a widely used paradigm for modeling dynamic decision-making, MSP finds numerous applications in addressing real-world control problems. There exists a substantial body of literature on MSP, and multiple recent works dedicating effort to consider a distributionally robust variant of it; see, for example, Shapiro et al. (2021, Chapter 7) and Huang et al. (2017); Shapiro (2021); Pichler and Shapiro (2021). Consequently, in this section, we delve into generalizations of the MSP formulation of DRMDP. We aim to achieve a formulation equivalent to scenarios where a controller with an arbitrary action set contends with a history-dependent S-rectangular adversary with an arbitrary action set. Additionally, we establish connections between our formulations and these generalized MSPs by leveraging the corresponding DPPs.
We now study the finite horizon problem to avoid the issue of needing to make sense of countably many nested stochastic programs. We note that an infinite horizon version could be defined through a limiting argument, due to the presence of the discount. However, introducing this will distract the main message in this section. We focus on the case of a history-dependent controller and adversarial policy classes and explain how the max-min control problem defined by (2.16) can be seen as a variant of the MSP.
Before introducing the multistage stochastic programs, we define some notations. let $T+1$ denote the time horizon. We consider a general controller action distribution set $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ and adversarial action set $\mathcal{P}\subset\left\{{S\times A\rightarrow\mathcal{P}(\mathcal{S})}\right\}$.
The SA-rectangular marginalization of $\mathcal{P}$ (see (2.11)) is defined as $\overline{\mathcal{P}}_{s,a}:=\left\{{p_{s,a}:p\in\mathcal{P}}\right\}$ for all $s\in S$. We consider the product action set:
$\overline{\mathcal{P}}^{\mathrm{SA}}:=\prod_{(s,a)\in S\times A}\overline{\mathcal{P}}_{s,a}=\left\{{p=\left\{{p_{s,a}:(s,a)\in S\times A}\right\}:p_{s,a}\in\overline{\mathcal{P}}_{s,a}}\right\}.$
Also, the S-rectangular marginalization of $\mathcal{P}$ is defined as $\overline{\mathcal{P}}_{s}:=\left\{{p_{s}:p\in\mathcal{P}}\right\}$ for all $s\in S$. We also consider the product action set: $\overline{\mathcal{P}}^{\mathrm{S}}:=\prod_{s\in S}\overline{\mathcal{P}}_{s}$.
To simplify notation, we recall the bracket integral notation in (2.13). We further identify that for function $f:h_{T+1}\rightarrow\mathbb{R}$,
$\mu[f(h_{T+1})]$ $=\sum_{s_{0}\in S}\mu[s_{0}]f(s_{0},\dots,a_{T},s_{T+1}),$
$d_{t}(h_{t})[f(h_{T+1})]$ $=\sum_{a_{t}\in A}d_{t}(h_{t})[a_{t}]f(s_{0},\dots,s_{t},a_{t},\dots,s_{T+1}),$
$\kappa_{t}(g_{t})[f(h_{T+1})]$ $=\sum_{s_{t}\in S}\kappa_{t}(g_{t})[s_{t+1}]f(s_{0},\dots,a_{t},s_{t+1},\dots,s_{T+1}).$
The style of this section is a brief exposition rather than a detailed discussion. One can find a more refined discussion of the MSP formulations and equivalent marginalization of general rectangular action sets in Li and Shapiro (2023).
6.1.1 SA-Rectangular Formulation
Let $u^{\mathrm{SA}}$ be the solution associated with the DR Bellman equation (3.1) with $\mathcal{P}_{s}=\prod_{a\in A}\overline{\mathcal{P}}_{s,a}$. We use $r_{T+1}(s,a)=u^{\mathrm{SA}}(s)$ as the terminal reward. We first consider the following SA-rectangular MSP formulation:
$v_{\mathrm{H}}^{\mathrm{SA}}(\mu,\mathcal{Q},\mathcal{P},u^{\mathrm{SA}})$ $:=\mu\left[\sup_{d_{0}(h_{0})\in\mathcal{Q}}d_{0}(h_{0})\Bigg{[}r(s_{0},a_{0})+\right.$ (6.1)
$\left.\left.\gamma\inf_{\kappa_{0}(\cdot,\cdot)\in\mathcal{P}}\kappa_{0}(g_{0})\left[\sup_{d_{1}(h_{1})\in\mathcal{Q}}d_{1}(h_{1})\left[\dots\sup_{d_{T}(h_{T})\in\mathcal{Q}}d_{T-1}(h_{T-1})\left[\Xi^{\mathrm{SA}}_{T-1}\right]\right]\right]\right]\right]$
where
$\Xi^{\mathrm{SA}}_{T-1}:=$
$r(s_{T-1},a_{T-1})+\gamma\inf_{\kappa_{T-1}(g_{T-2},\cdot,\cdot)\in\mathcal{P}}\kappa_{T-1}(g_{T-1})\left[\sup_{d_{T}(h_{t})\in\mathcal{Q}}d_{T}(h_{T})\left[r(s_{T},a_{T})+\gamma\inf_{\kappa_{T}(g_{T},\cdot,\cdot)\in\mathcal{P}}\kappa_{T}(g_{T})[u^{\mathrm{SA}}]\right]\right].$
We claim that
$v_{\mathrm{H}}^{\mathrm{SA}}(\mu,\mathcal{Q},\mathcal{P},u^{\mathrm{SA}})=v_{\mathrm{H}}^{\mathrm{SA}}(\mu,\mathcal{Q},\overline{\mathcal{P}}^{\mathrm{SA}},u^{\mathrm{SA}}).$
To see this, let’s consider the first $\inf$ taken over $\kappa_{0}(\cdot,\cdot)\in\mathcal{P}$. Since it is inside the second square bracket, where the sum over $s_{0}$ and $a_{0}$ is taken outside, the inf is taken for each history $g_{0}=(s_{0},a_{0})$ separately. In particular, the adversary can choose different $\kappa_{0}$ different $(s_{0},a_{0})$. Thus, the inf over $\kappa_{0}(\cdot,\cdot)\in\mathcal{P}$ is equivalent to choosing from $\overline{\mathcal{P}}^{\mathrm{SA}}$. For $t>0$, the same holds true analogously. This observation implies the claim.
Upon writing the dynamic programming equations for this program, one realizes that
$v_{\mathrm{H}}^{\mathrm{SA}}(\mu,\mathcal{Q},\mathcal{P},u^{\mathrm{SA}})=v_{\mathrm{H}}^{\mathrm{SA}}(\mu,\mathcal{Q},\overline{\mathcal{P}}^{\mathrm{SA}},u^{\mathrm{SA}})=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}})$
where $\Pi^{\mathrm{C}}_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}^{\mathrm{SA}}$ has action sets $\mathcal{Q}$ and $\overline{\mathcal{P}}_{s,a}:(s,a)\in S\times A$ respectively.
We remark that in their endeavor to formulate a MSP with adversarial dynamics, Huang et al. (2017) successfully derived a formulation that is a special class of instances of the MSP formulation presented in this section. This consequently results in a DPP that is equivalent to the SA-rectangular case detailed in this paper.
6.1.2 S-Rectangular Formulation
Based on the same core idea, we can also define a S-rectangular MSP from the general adversarial action set $\mathcal{P}$. Concretely, let $u^{\mathrm{S}}$ be the solution to (3.1) with $\mathcal{P}_{s}=\overline{\mathcal{P}}_{s}$, we consider
$v_{\mathrm{H}}^{\mathrm{S}}(\mu,\mathcal{Q},\mathcal{P},u^{\mathrm{S}}):=\mu\left[\sup_{d_{0}(h_{0})\in\mathcal{Q}}\inf_{\kappa_{0}(\cdot,\cdot)\in\mathcal{P}}d_{0}(h_{0})\left[r(s_{0},a_{0})+\gamma\kappa_{0}(g_{0})\left[\sup_{d_{1}(h_{1})\in\mathcal{Q}}\inf_{\kappa_{1}(g_{0},\cdot,\cdot)\in\mathcal{P}}d_{1}(h_{1})\left[\dots\right]\right]\right]\right]$ (6.2)
with terminal reward $r_{T+1}(s,a)=u^{\mathrm{S}}(s)$, where the final terms are handled in a similar fashion as $\Xi^{\mathrm{S}A}_{T-1}$, except with the infimum taken outside $d$. Similar to the SA-rectangular version, we have that
$v_{\mathrm{H}}^{\mathrm{S}}(\mu,\mathcal{Q},\mathcal{P},u^{\mathrm{S}})=v_{\mathrm{H}}^{\mathrm{S}}(\mu,\mathcal{Q},\overline{\mathcal{P}}^{\mathrm{S}},u^{\mathrm{S}})=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ (6.3)
where $\Pi^{\mathrm{C}}_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ has action sets $\mathcal{Q}$ and $\overline{\mathcal{P}}_{s}:s\in S$ respectively. The first equality can be seen by a similar reason. Let us consider the second infimum taken over $\kappa_{1}(g_{0},\cdot,\cdot)\in\mathcal{P}$ as an example. The $\inf$ is within the sum over $s_{0},a_{0},s_{1}$. In particular, the adversary can choose different $\kappa_{1}$ for each $s_{1}$ observed. Therefore, one can see that the use of $\mathcal{P}$ for the MSP problem $v_{\mathrm{H}}^{\mathrm{S}}(\mu,\mathcal{Q},\mathcal{P})$ effectively gives the adversary the freedom to choose from the larger set $\overline{\mathcal{P}}^{\mathrm{S}}$. This observation is consistent with the behavior of formulation (2.1) in Li and Shapiro (2023). In their work, (2.1) formulates the transition dynamics as a two-player zero-sum game. In this setting, they also establish an equivalence relationship similar to (6.3).
6.1.3 General Rectangular Case: Some Comments
In the same spirit, one can also define a generalization of the MSP with terminal reward $u$ in which the adversary cannot observe the realization of the current state, leading to the adversary’s need to choose its action from the general action set. However, caution is required to maintain the sup-inf structure. This is because the controller is one who can use the realization of the current state as available information, while the adversary is not, leading to the temptation to consider a formulation that puts the inf over $\kappa_{t}$ outside the sum of $\kappa_{t-1}$. However, doing so directly will result in the reverse of the ordering of the inf and sup. To address this, we can redefine the controller as one who selects, for each and every state, an action distribution; i.e., $d_{t}(g_{t-1},\cdot)\in\left\{{S\rightarrow\mathcal{Q}}\right\}$. We note that the MSP for both the SA and S-rectangular cases in the preceding sections can also adopt this equivalent formulation of the controller. Then, the sup can be moved outside of the sum with respect to the preceding $\kappa$. We didn’t adopt this reformulation of the controller in the SA and S-rectangular cases so as to keep it consistent with the DRMDP formulations. This leads to the following definition of a generalized MSP:
$v_{\mathrm{H}}^{\mathrm{N}}(\mu,\mathcal{Q},\mathcal{P},u):=$ (6.4)
$\sup_{d_{0}(\cdot)\in\left\{{S\rightarrow\mathcal{Q}}\right\}}\inf_{\kappa_{0}(\cdot,\cdot)\in\mathcal{P}}\mu\left[d_{0}(h_{0})\left[r(s_{0},a_{0})+\gamma\sup_{d_{1}(g_{0},\cdot)\in\left\{{S\rightarrow\mathcal{Q}}\right\}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\inf_{\kappa_{1}(g_{0},\cdot,\cdot)\in\mathcal{P}}\kappa_{0}(g_{0})}\left[d_{1}(h_{1})\left[\dots[u]\right]\right]\right]\right],$
for some terminal values $u$. Observe that this is not a MSP in the classical sense: the segment highlighted in blue shows that the inf over $\kappa_{1}$ is between inf over $\kappa_{0}$ and the expectation w.r.t. $\kappa_{0}$.
Nevertheless, one faces difficulties when trying to conclude that the max-min control value in (2.16) with the general adversarial policy class $\mathrm{K}_{\mathrm{H}}^{\mathrm{N}}$ and controller’s policy class $\Pi^{\mathrm{C}}_{\mathrm{H}}$ is equivalent to this generalized MSP. This is attributed to the absence of a DPP—as mentioned earlier, the max-min control value in this context doesn’t satisfy a Bellman equation in full generality. One can envision a similar situation for the generalized MSP value. Notably, the $\inf$ over $\kappa_{t}$ is always positioned outside of the expectation over $\kappa_{t-1}$, as exemplified in blue in the above definition (6.4). This suggests that determining the optimal $\kappa_{t}$ requires considering the choice of $\kappa_{t-1}$, taking into account both the history and the preceding choices. In turn, $\kappa_{t-1}$ depends on the selection of $\kappa_{t-2}$, and so on, resulting in a fully intertwined relationship. This intricate interplay differs from the SA-rectangular (6.1) and S-rectangular (6.2) formulations and it falls beyond the scope of this paper.
In summary, it is possible to formulate a robust dynamic decision-making environment, namely the generalized MSP $v_{\mathrm{H}}^{\mathrm{N}}(\mu,\mathcal{Q},\mathcal{P},u)$, in which an adversary chooses from a general rectangular ambiguity set $\mathcal{P}$ possibly yields different value as one that chooses from $\overline{\mathcal{P}}^{\mathrm{S}}$; i.e.
$v_{\mathrm{H}}^{\mathrm{N}}(\mu,\mathcal{Q},\overline{\mathcal{P}}^{\mathrm{S}},u)\leq v_{\mathrm{H}}^{\mathrm{N}}(\mu,\mathcal{Q},\mathcal{P},u),$
which contrasts with the equality (6.3). Notably, instances can be formulated where the aforementioned inequality changes to strict inequality ($<$), even when $\mathcal{Q}=\mathcal{P}(\mathcal{A})$. As an illustrative example, consider the inventory control setting in Example 1 with an adversary that lacks visibility of the current inventory level $I_{t}$ (general rectangular) compared to one that has full information about $I_{t}$ available for decision-making (S-rectangular).
In addition, the connection between this generalized MSP and DRMDP with history-dependent controllers and general-rectangular adversaries remains an open research question; i.e., whether there exists a terminal reward $u$ s.t.
$v_{\mathrm{H}}^{\mathrm{N}}(\mu,\mathcal{Q},\mathcal{P},u)\stackrel{{\scriptstyle?}}{{=}}v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{N}})$
holds all $\mu\in\mathcal{P}(\mathcal{S})$. It remains an intriguing research avenue that extends beyond the scope of the present paper.

### v4

@@ 5.3 Deterministic Controllers
Finally, we analyze the case with a history-dependent deterministic controller and a Markov convex adversary. We consider the history-dependent policy such that $\pi_{2i}=a_{1},i=1,2,\ldots,$ if the controller sees state $\text{I}_{1}$ and $\pi_{2i}=a_{2},i=1,2,\ldots,$ if the controller sees state $\text{I}_{2}$. Then, for any Markov adversary $\kappa$, we always have $v(\delta_{\text{I}_{0}},\pi,\kappa)=0>u^{*}(\text{I}_{0}).$ Again, this suggests the absence of a DPP in this setting as well.
6 Asymptotically Optimal History-Dependent Control
As demonstrated in Section 5.1, the DPP does not generally hold when a history-dependent controller interacts with a time-homogeneous adversary selecting actions from a non-convex S-rectangular set. This breakdown implies that Markov time-homogeneous policies are strictly suboptimal and that an effective policy in this setting must intelligently adapt to the historical information it generates. As previously discussed, this scenario is of practical significance, as real-world decision-making often allows the flexibility to employ non-stationary strategies in a potentially adversarial environment that remains completely static. Given these considerations, a natural question arises: Can we design good history-dependent policies that are both theoretically sound and implementable in real-world applications?
In this section, we provide an affirmative answer by proposing a reinforcement learning (RL)-based history-dependent policy that can be effectively implemented in various RMDP settings. We demonstrate that this policy is asymptotically optimal as the effective horizon $1/(1-\gamma)$ approaches infinity, as formally defined in Definition 8. Crucially, our proposed policy is built on the intuition developed in Section 5: to achieve asymptotic optimality, the controller must exploit the adversary’s time-homogeneous nature while effectively learning the fixed transition structure imposed by the adversary. Moreover, this strategic adaptation must ensure sufficient exploration and the near-optimal utilization of historical data. Together, these elements enable the policy to achieve near-optimal performance over long horizons.
Concretely, we analyze a scenario in which a time-homogeneous adversary interacts with a history-dependent controller. For clarity in this presentation, we assume $\mathcal{Q}=\mathcal{P}(\mathcal{A})$; however, we will later discuss in Remark 6 how the main result can be extended to non-convex $\mathcal{Q}$. We aim to demonstrate that, given an adversarial policy class $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ with non-convex S-rectangular ambiguity sets $\{\mathcal{P}_{s}:s\in S\}$, we can construct an asymptotically optimal history-dependent policy $\pi\in\Pi_{\mathrm{H}}$ in the following sense.
Definition 8.
We say that a history-dependent policy $\pi\in\Pi_{\mathrm{H}}$ (which may depend on $\gamma$) is asymptotically optimal for the RMDP if
$0\leq(1-\gamma)[v_{\gamma}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})-v_{\gamma}(\mu,\pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})]\rightarrow 0$ (6.1)
as $\gamma\rightarrow 1$. Note that here we emphasize the $\gamma$ dependence of the value function by adding a subscript.
Remark 5.
If we further assume that the controller can achieve positive reward at any state; i.e. $r_{\vee}(s):=\max_{b\in A}r(s,b)>0$ for all $s\in S$, then $v_{\gamma}(s,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\geq\frac{1}{1-\gamma}r_{\vee}(s)$. Therefore, (6.1) implies that
$1\geq\frac{v_{\gamma}(\mu,\pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})}{v_{\gamma}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})}$ $=1-\frac{v_{\gamma}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})-v_{\gamma}(\mu,\pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})}{v_{\gamma}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})}$
$\geq 1-\frac{(1-\gamma)[v_{\gamma}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})-v_{\gamma}(\mu,\pi,\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})]}{E_{\mu}[r_{\vee}(X_{0})]}\rightarrow 1$
i.e. the approximation ratio goes to $1$ as $\gamma\rightarrow 1$.
Building on the intuition that the controller can potentially exploit the time-homogeneous adversary’s inability to adapt and learn the full dynamics of the adversary, we propose the following Policy 1 as a candidate for asymptotic optimality in the sense of Definition 8.
Input: Discount factor $\gamma\in(0,1)$ and exploration period $n\geq 1$.
Let $\bar{\pi}_{0}\dots\bar{\pi}_{n}$ and $\hat{p}$ be the output of the exploration Subroutine 2.
[…]
Policy 1 Explore-then-Exploit Policy
Policy 1 has two phases. During the first phase, the policy use the decisions generated by the Exploration Subroutine 2 that explores the entire state and action spaces and gathers a data set. Using these data, the policy constructs a deterministic decision rule as described in the Exploitation Subroutine 3 and uses it thereafter. We show that under the assumption that the all $p\in\mathcal{P}^{S}$ is communicating [55, 88], then this algorithm is asymptotically optimal in the above sense.
Input: Exploration period $n\geq 1$.
Choose the effective sample size
$m=\left\lfloor{\frac{n}{8|Z|D^{3}|A|^{D}}}\right\rfloor$ (6.2)
and define random variables $\tau_{1,1}=\inf_{t\geq 0}\left\{{s_{t}=z^{1}(0)}\right\}$ and for $k=1,2,\dots,m$,
$\tau_{j,k+1}:=\inf_{t\geq\tau_{j,k}+1}\left\{{s_{t}=z^{j}(0)}\right\},\quad\tau_{j+1,1}:=\inf_{t\geq\tau_{j,m}+1}\left\{{s_{t}=z^{j+1}(0)}\right\}.$ (6.3)
This defines a sequence of $\mathcal{H}_{t}$ stopping times of consecutively hitting each of $s\in S$ for $|A|m$ times.
Since $\omega\rightarrow\tau_{j,k}(\omega)$ is $\mathcal{H}_{t}$ measureable, we can define $B_{j,k,t}(h_{t}):=\mathds{1}\left\{{\tau_{j,k}(\omega)\leq t,\omega=(h_{t},\dots)}\right\}$ and $K_{j,t}(h_{t}):=\sum_{k=1}^{m}B_{j,k,t}(h_{t})$, and set
$J_{t}(h_{t})$ $=1+\sum_{j=1}^{|Z|}\mathds{1}\left\{{K_{j}(h_{t})=|A|m}\right\}$
$B_{t}(h_{t})$ $=B_{J_{t},K_{J_{t},t}+1}(h_{t})\mathds{1}\left\{{J_{t}(h_{t})\leq|Z|}\right\}.$
Naturally, these definitions extend to random variables with $B_{j,k,t}(\omega)=B_{j,k,t}(H_{t})$.
Output 1: Construct the sequence of decision rules for the exploration phase by defining
$\bar{\pi}_{t}(h_{t})(a)=B_{t}(h_{t})\mathds{1}\left\{{z^{J}(1)=a}\right\}+(1-B_{t}(h_{t}))\frac{1}{|A|}$
for all $h_{t}\in\mathbf{H}_{t}$ and $a\in A$. Return $\bar{\pi}_{0},\dots,\bar{\pi}_{n}$.
Output 2: Recall the state projection mapping $X_{t}(\omega)=s_{t}$ where $\omega=(s_{0},a_{0},\dots,s_{t},\dots)$. We let $X_{z^{j}}^{(k)}=X_{\tau_{j,k}+1}$, and define the following mixture kernel $\hat{p}=\left\{{\hat{p}_{z}\in\mathcal{P}(\mathcal{S}):z\in Z}\right\}$, seen as a random element on $(\Omega,\mathcal{H}_{n})$, where
$\hat{p}_{z}(s^{\prime})=\mathds{1}\left\{{J_{n}=|Z|+1}\right\}\frac{1}{m}\sum_{k=1}^{m}\mathds{1}\left\{{X_{z^{j}}^{(k)}=s^{\prime}}\right\}+\mathds{1}\left\{{J_{n}\leq|Z|,s^{\prime}=z(0)}\right\}.$
Return $\hat{p}$.
Subroutine 2 Exploration
To introduce the Exploration and Exploitation subroutines, we define some new notations. For notation simplicity, we index every $(s,a)\in S\times A$ by denoting $S\times A:=|Z|=\left\{{z^{1},\dots,z^{|Z|}}\right\}$. Also, for $z=(s,a)$, we write $z(0)=s$ and $z(1)=a$. Let $\mathcal{H}_{t}=\sigma(H_{t})$ be the $\sigma$-field generated by $H_{t}=(X_{0},A_{0},\dots A_{t-1},X_{t})$.
Input: Exploration period $n\geq 1$, discount factor $\gamma\in(0,1)$, and empirical kernel $\hat{p}$ seen as a measurable function on $(\Omega,\mathcal{H}_{n})$.
Conditioned on $\mathcal{H}_{n}$, solve for the unique fixed point of the Bellman equation
$\hat{v}(s)=\max_{a\in A}r(s,a)+\gamma\sum_{s^{\prime}\in S}\hat{p}_{s,a}(s^{\prime})\hat{v}(s^{\prime}),\quad\forall s\in S.$
Note that $\hat{v}\in m\mathcal{H}_{n}$.
Output: For all $t\geq n+1$, $h_{t}\in\mathbf{H}_{t}$, and $a\in A$, set
$\hat{\pi}_{t}(a|h_{t})=\frac{\mathds{1}\left\{{a\in\arg\max_{b\in A}\hat{r}(s,b)+\gamma\sum_{s^{\prime}\in S}\hat{p}_{s,b}(s^{\prime})\hat{v}(s^{\prime})}\right\}}{\left|\arg\max_{b\in A}\hat{r}(s,b)+\gamma\sum_{s^{\prime}\in S}\hat{p}_{s,b}(s^{\prime})\hat{v}(s^{\prime})\right|}.$
Return $\hat{\pi}$.
Subroutine 3 Exploitation—Empirical Value Iteration (EVI)
We provide an intuitive description of Policy 1. The sequence of decisions $\bar{\pi}_{0},\dots,\bar{\pi}_{n},\dots$ generated by the Exploration Subroutine 2 achieves the following mechanism for any given integer $m\geq 1$.
  1. 1.
    Starting from an initial state, use a uniformly random decision until hitting $s^{1}$. Set $j=1$, $k=0$, $l=1$.
  2. 2.
    Use action $a^{l}$ to transition to a new state and set $k\leftarrow k+1$. If $k=m$, and $l\leq|A|-1$, we move on to the next action by assigning $l\leftarrow l+1$. Else, if $k=m$, and $l=|A|$, moving onto the next state by assigning $j\leftarrow j+1$ and starting from action $l=1$.
  3. 3.
    Use the uniformly random decision until hitting state $s^{j}$. In particular, if $j>|S|$, the uniformly random decision is used indefinitely. Upon hitting $s^{j}$, go back to step 2.
Under the assumption that every $p\in\mathcal{P}^{S}$ is communicating and $n$ is sufficiently large, this subroutine can, with high probability, successfully collect $m$ i.i.d. transition samples starting from each $z\in Z$, sampled from the adversarial transition kernel fixed over time and history. Upon the successful collection of the transition samples, we form an empirical transition kernel $\hat{p}$.
The exploitation phase, as outlined in Subroutine 3, simply applies the optimal policy obtained from the empirical transition kernel $\hat{p}$ by solving the corresponding non-robust empirical Bellman equation. We note that the policy in Subroutine 3 randomizes the action in cases of tied values. Alternatively, one could opt for deterministically selecting one action when ties occur.
In summary, under Policy 1, the controlled Markov chain will, with high probability, successfully collect an empirical kernel consisting of $m$ i.i.d. transition samples for each state-action pair. Subsequently, an empirical optimal policy is computed using the empirical value iteration subroutine, which, with high probability, yields a policy that is close to optimal. Concretely, Policy 1 satisfies the following Theorem:
Theorem 6 (Asymptotic Optimality of Policy 1).
Assume that for every $\in\mathcal{P}^{\mathrm{S}}, pp$ is communicating with diameter at most $D$ (c.f. Zurek and Chen, [88]). Then, Policy 1 with $n=\frac{1}{\sqrt{1-\gamma}}$ is asymptotically optimal in the sense of (6.1).
Remark 6.
One can extend this result to establish the asymptotic optimality of policies analogous to Policy 1 in the deterministic controller setting $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. However, weak communication alone may not suffice, as uniform randomization cannot be leveraged for exploration. Instead, we can assume the existence of a Markov time-homogeneous decision rule $\bar{\pi}$ such that, for all $p\in\mathcal{P}$, the transition kernel $(s,s^{\prime})\mapsto p_{s,\bar{\pi}(s)}(s^{\prime})$ induces an irreducible Markov chain with a bounded mixing time $t_{\text{mix}}$.
Under this condition, substituting $\bar{\pi}$ for the uniform randomization policy $1/|A|$ in the Exploration Subroutine 2, along with an appropriately adjusted choice of $m$, ensures that a version of Theorem 6 holds in the setting where $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$.
To conclude this section, we note that readers familiar with the sample complexity literature in RL might, based on the intuition established earlier, consider model-free algorithms such as asynchronous Q-learning [78] as potential candidates for asymptotically optimal algorithms. However, due to its non-minimax optimal sample complexity dependence on $1/(1-\gamma)$ [36], Q-learning does not yield an asymptotically optimal policy in this setting. Instead, achieving asymptotic optimality requires more refined model-free algorithms that are optimal under appropriate mixing conditions, such as the variance-reduced Q-learning method proposed in [75].
@@ 7 Connections with other Formulations
In this section, we elucidate the links between the formulations of RMDP and the multistage stochastic programs [61, Chapter 3] within the realm of optimization literature. Additionally, we explore their connections with stochastic games, as introduced by Shapley, [64] in the economics literature.
@@ 7.1 Multistage Stochastic Programs
Early works on RMDP, such as González-Trejo et al., [18] and Nilim and El Ghaoui, [47], interpreted the max-min problem as dynamic interactions between a controller and an SA-rectangular adversary. Indeed, the presence of a Dynamic Programming Principle (DPP) in these scenarios implies that the interactions under optimal robust control are equivalent to another dynamic model wherein the controller and adversary sequentially make decisions. This interpretation results in what is now considered as a robustified variant of the multistage stochastic program (MSP) [27]. In contrast, Nilim and El Ghaoui, [47]’s version of sequential min-max-…-min-max decisions (see Equation (5) therein) is selected before the state and action sequences are realized. As a widely used paradigm for modeling dynamic decision-making, MSP finds numerous applications in addressing real-world control problems. There exists a substantial body of literature on MSP, and multiple recent works dedicating effort to consider a robust variant of it; see, for example, Shapiro et al., [61, Chapter 7] and Huang et al., [27], Shapiro, [59], Pichler and Shapiro, [53].